A cold water tap can fill a tub in 6 minutes and a hot water tap can fill the tub in 8 minutes. A drain can empty a full tub in 10 minutes. If both taps are on and the drain is open, how long will it take to fill the tub?
Answer by lwsshak3(6456)

A cold water tap can fill a tub in 6 minutes and a hot water tap can fill the tub in 8 minutes. A drain can empty a full tub in 10 minutes. If both taps are on and the drain is open, how long will it take to fill the tub?
**
1/6 =work rate of cold water tap
1/8=work rate of hot water tap
1/10=work rate of drain
let x=minutes to fill tub when both taps and drain are open
..
x/6+x/8-x/10=100%=1
LCD:120
20x+15x-12x=120
23x=120
x=120/23 ≈5.22 minutes
ans:
Time it takes to fill tub when both taps and drain are open=5.22 minutes